Evaluate 1/2(1.5+3/4÷1/4)
Found 2 solutions by math_helper, greenestamps:
Answer by math_helper(2461)

1/2(1.5+3/4÷1/4)

By PEMDAS, this is the same as (1/2)*(1.5+3/4+1/4)
which is (1/2)*(2.5) or 1.25 (or 5/4 as a fraction)
-----
@tutor greenestamps: it is 1/4 inside the paren's not 4/1...
Never mind, I see now it is a divide symbol, not a plus + sign :-(
1/2(1.5+3/4 / (1/4)) = (1/2)(1.5+(3/4)*(4/1)) = (1/2)(4.5) = 2.25, yes... sry

Answer by greenestamps(13200)



PEMDAS:
P: simplify inside parentheses first (there are parentheses in this expression)
E: do exponents next (there are none here)
MD: next do multiplication and division, in order left to right
AS: last do addition and subtraction, again in order left to right

Inside the parentheses, there is an addition followed by a division. Do the division first.

1/2(1.5+3/4÷1/4) --> 1/2(1.5+(3/4)(4/1)) = 1/2(1.5+3)

Continue simplifying inside the parentheses.

1/2(1.5+3) --> 1/2(4.5)

Last do the multiplication.

1/2(4.5) = 2.25